OpenStudy (anonymous):

How to find four geometric means between 243 and 1?

OpenStudy (anonymous):

the geometric mean between a and b is sqrt(ab)

OpenStudy (anonymous):

sqrt(ab)?

OpenStudy (anonymous):

You can treat 243 as \(a_1\) of a geometric series and 1 as \(a_5\) and use the formula for the \(n^{th}\) term of a geometric series to find r.

OpenStudy (anonymous):

I dont know the equation for geometric series..

OpenStudy (anonymous):

\(a_n = a_1 * r^{n-1}\)

OpenStudy (anonymous):

\(a_5 = 1 = a_1 * r^{5-1} = 243 * r^4\)

OpenStudy (anonymous):

So r is equal to the fourth root of \(\frac{1}{243}\)

OpenStudy (anonymous):

To get 4 means in between you will need to set 1 = a_6 instead though, sorry.

OpenStudy (anonymous):

Which will mean the fifth root instead of the fourth root.

OpenStudy (anonymous):

a5= 1 and a1= 1?

OpenStudy (anonymous):

\[a_n = 243 * \left(\sqrt[5]{\frac{1}{243}}\right)^{n-1}\]

OpenStudy (anonymous):

It should be \(a_1\) = 243 and \(a_6=1\).

OpenStudy (anonymous):

how do we know it is the 6th term?

OpenStudy (anonymous):

243 is the first term, and you want 4 terms to between 243 and one. So the last term you're trying to find will be the fifth one, and 1 is then the 6th term.

OpenStudy (anonymous):

Like this: 243, \(a_2\), \(a_3\), \(a_4\), \(a_5\), 1

OpenStudy (anonymous):

so i do \[a _{5}=243(r ^{5-1})\] right?

OpenStudy (anonymous):

You could do that to find the fifth term, but because you know the sixth term in order to find r you'll want to use 6 instead of 5.

OpenStudy (anonymous):

oh! yeah..

OpenStudy (anonymous):

Because you know \(a_6 = 1\) you can solve for r if you set \(n=6\).

OpenStudy (anonymous):

so why is it like you said earlier? the \[\frac{ 1 }{ 243 }\]

OpenStudy (anonymous):

i know it is \[a _{6}=243 \times (r ^{5})\] but i dont know how to solve it completely..

OpenStudy (anonymous):

When you set \(n=6\), you get \(a_6 = a_1 * r^{n-1}\), then \( 1 = 243 * r^5 \).

OpenStudy (anonymous):

where did the 1 come from?

OpenStudy (anonymous):

oh nevermind!

OpenStudy (anonymous):

You know \(a_6 = 1\).

OpenStudy (anonymous):

okay so now how do i finish solving it?

OpenStudy (anonymous):

wait i get it.. but what does it come out to if i dont have a calculator?

OpenStudy (anonymous):

\[r=\sqrt[5]{\frac{ 1 }{ 243 }}\] is what i got it down to

OpenStudy (anonymous):

\(\sqrt[5]{243} = 3\) so you can simplify it if you split the fraction up.

OpenStudy (anonymous):

how did you split it up? ive never done that

OpenStudy (anonymous):

and where did r go?

OpenStudy (anonymous):

That was just a general statement, not part of the equation.

OpenStudy (anonymous):

I was just saying that you can simplify that root down to a nice number.

OpenStudy (anonymous):

yes but how did you do that

OpenStudy (anonymous):

You could then say that \(r = \frac{1}{3}\).

OpenStudy (anonymous):

\(\sqrt{ \frac{a}{b} } = \frac{\sqrt{a}}{\sqrt{b}}\) when a and b are both positive.

OpenStudy (anonymous):

ok but if i kept just what i had without simplifiying it down how would i do it?

OpenStudy (anonymous):

The easiest way would just be to leave everything in root form in the answer.

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

i.e. \(a_2 = 243 * \sqrt[5]{\frac{1}{243}}\)

OpenStudy (anonymous):

How much you're expected to simplify would really depend on the class and what unit you're doing etc. so it's hard to say exactly how far your particular teacher will want it simplified.

OpenStudy (anonymous):

do you know the formula for the sum of geometric series? i cant find it

OpenStudy (anonymous):

\[a_1 \frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

thank you